A man's age is three times his son's age. Ten years ago he was five times his son's age. By forming two equations and solving them simultaneously, find both of their ages.
Answer by ankor@dixie-net.com(22740)

A man's age is three times his son's age. Ten years ago he was five times his son's age. By forming two equations and solving them simultaneously, find both of their ages.
:
Write an equation for each statement
:
"A man's age is three times his son's age."
m = 3s
or
m - 3s = 0
:
" Ten years ago he was five times his son's age."
m - 10 = 5(s-10)
m - 10 = 5s - 50
m - 5s = =-50 + 10
m - 5s = -40
:
solving them simultaneously
m - 3s = 0
m - 5s = -40
---------------Subtraction eliminates m, find s
0 + 2s = 40
s = 20 yrs is the son's age
I'll let you find the man's age